# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v2

@@ Appendix F Details of the no-go theorems for 1D decoding of dynamical codes
Since there are $n/2$ measurements in a given time step, there can be at most $n/2$ independent detectors formed from such measurements. However, there will in fact be far fewer for a GDC, because each individual measurement anti-commutes with a measurement from the previous time step resulting in a random value for each edge measurement. Therefore, a detector with a deterministic outcome must be formed from at least a product of two non-deterministic measurements within a time step. Suppose the GDC has a $Z$-detector graph (edges in the ZDG correspond to particular $Z$ errors). Again, we consider the subgraph corresponding to $Z$ errors occurring in time steps $k_{i}<t<T-k_{f}$, where $T$ is the GDC’s duration, and detectors formed in time steps $t>k_{i}$. There are $<n(T-k_{i})/2$ vertices and $n(T-1-k_{i}-k_{f})$ edges in this subgraph. Since the number of vertices of a collection of repetition code decoding graphs cannot be less than the number of edges, this means that the ZDG of a GDC cannot be a collection of repetition code graphs for $T\geq 2k_{f}+k_{i}+2$. ∎

### v3

@@ IV.6 No-go theorems for 1D decoding graphs of dynamical codes
Since there are $n/2$ measurements in a given time step, there can be at most $n/2$ independent detectors formed from such measurements. However, there will in fact be far fewer for a GDC, because each individual measurement anti-commutes with a measurement from the previous time step resulting in a random value for each edge measurement. Therefore, a detector with a deterministic outcome must be formed from at least a product of two non-deterministic measurements within a time step. Suppose the GDC has a $Z$-detector graph (edges in the ZDG correspond to particular $Z$ errors). Again, we consider the subgraph corresponding to $Z$ errors occurring in time steps $t_{i}<t<T-t_{f}$, where $T$ is the GDC’s duration, and detectors formed in time steps $t>t_{i}$. There are $<n(T-t_{i})/2$ vertices and $n(T-1-t_{i}-t_{f})$ edges in this subgraph. Since the number of vertices of a collection of repetition code decoding graphs cannot be less than the number of edges, this means that the ZDG of a GDC cannot be a collection of repetition code graphs for $T\geq 2t_{f}+t_{i}+2$. ∎
Owing to this theorem, we do not expect to see thresholds in dynamical codes as high as 50%. The reason that static codes perform better than dynamical codes under biased noise is that they obey two separate symmetries under an infinitely phase-biased code-capacity noise model. There is a spatial symmetry that forces syndromes to appear in pairs along 1D strips of the lattice, but there is also a temporal symmetry: syndromes must appear in pairs within a given time slice. These two symmetries simplify the detector graph into a collection of disjoint repetition code decoding graphs, leading to a high performance (up to 50% threshold) of the static codes under biased noise. On the other hand, even for the case where measurement errors are not considered, GDCs necessarily break the temporal symmetry (for HDCs, this is a consequence of Lemma 1), leading to edges between vertices in different time slices of the detector graph (see Fig. 3). We thus expect that such codes can only possess a single symmetry under an infinitely phase-biased code-capacity noise model, as the X3Z3 Floquet code does.
Acknowledgements.
We would like to thank our colleagues at Riverlane, especially Alexandra Moylett, Ophelia Crawford, and György Gehér for helpful discussions. We are also grateful to Ophelia Crawford and Hari Krovi for their comments on the manuscript. CM would like to thank Benjamin Brown for pointing out a useful circuit for performing next-nearest-neighbour entangling gates. CM acknowledges support from the Intelligence Advanced Research Projects Activity (IARPA), under the Entangled Logical Qubits program through Cooperative Agreement Number W911NF-23-2-0223.
Author contributions
Data availability
The threshold data is available in https://doi.org/10.5281/zenodo.14258878 (Ref. [69]). Other numerical data generated in this work is available from the authors upon reasonable request. The video talk on this manuscript is available in https://www.youtube.com/watch?v=nqQT-5IRC9w.
Code availability
Circuits used in the simulations are available in https://doi.org/10.5281/zenodo.14258878 (Ref. [69]).
Competing interests
The authors declare no competing interests.